Circular No. 5973

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) For the Hubble Space Telescope (HST) observing team, H. A. Weaver,

Space Telescope Science Institute, reports the following results from their analysis of HST WFPC-2 images and FOS spectra taken during Mar. 29–30: “Fragment 8b is now clearly two separate fragments (provisionally designated 8b_{1} and 8b_{2} for the brighter and fainter ones, respectively). Fragment 8b_{2} appears rather diffuse and is a good candidate for disappearing over the next month or so. At the expected positions of fragments 4 and 8a, there are only barely discernible ‘puffs’, which will presumably become undetectable within a month. The brightness ‘spur’ near fragment 5 (cf. IAUC 5947) has apparently decreased in intensity, and there is no evidence for any new fragments in this region. All of the other fragments still have strongly condensed appearances, much as they had during the previous HST observations in July 1993 and Jan. 1994. The amount of dust within ∼ 1′′ of fragment 15 (the best-analyzed) has apparently changed very little since last July, while the amount of dust farther out decreased dramatically (e.g., a decrease by a factor of ∼ 4 from July to Jan. for a 5′′-radius aperture; there has been little change in the dust content within 5′′ from Jan. to Mar.). No emission features are seen in an 88-min FOS spectrum of the region 222–328 nm with the 3′′.7 × 1′′.2 aperture placed near fragment 7a. The limit on the brightness of the OH (0,0) band near 309 nm is 1.9 Rayleighs, which translates into an upper limit of ∼ 2.4 × 10^{11} molecules/cm^{2} for the average OH column density and ∼ 1.2 × 10^{27} molecules/s for the water production rate (all limits are 3σ).”

SUPERNOVA 1994I IN NGC 5194 P. Wild, Astronomical Institute, Berne, reports that a photograph on

T-Max film taken Mar. 30.916 UT at Zimmerwald (limiting magnitude near V = 16.0) tentatively suggests that the supernova was not visible, although the bright galaxy background makes it difficult to draw a definite conclusion.

SUPERNOVA 1994K IN ANONYMOUS GALAXY Corrigendum. On IAUC 5972, line 8, for (University of Kentucky) read

(Multiple Mirror Telescope Observatory). Also, the spectrogram reported by Schmidt and Kirshner was taken on Apr. 10 UT.

1994 April 12 Daniel W. E. Green